# When Tokens Talk Too Much: A Survey of Multimodal Long-Context Token Compression across Images, Videos, and Audios

### v4

@@ III-D2 Cross-Modal Selection
FastV [26] 50% 47.95 3.47 3.36 3.77 3.50 2.99 2.57 81.1 58.0 57.5
DyCoke [17] 50% 47.88 3.47 3.33 3.76 3.51 3.01 2.58 81.1 57.7 57.4 57.5
PLLaVA [3] 50% 47.59 3.45 3.36 3.73 3.52 3.00 2.66 81.0 57.7 56.9
LLaVA-Scissor [102] 50% 47.89 3.47 3.37 3.76 3.47 3.00 2.65 81.12 57.6 57.4
@@
FastV [26] 35% 47.83 3.46 3.32 3.74 3.47 2.97 2.61 80.5 57.8 56.0 61.3
DyCoke [17] 35% 47.81 3.45 3.31 3.74 3.46 2.98 2.54 80.9 57.7 56.2 61.8
PLLaVA [3] 35% 47.23 3.42 3.26 3.70 3.39 2.92 2.59 79.66 56.07 54.26 59.5
VisionZip [18] 25% 60.3 56.5 58.2 57.9
PruneVid [27] 25% 59.9 55.7 57.4 57.4
FastVID [44] 25% 56.3 58.0 56.5
LLaVA-Scissor [102] 25% 47.79 3.47 3.33 3.76 3.47 2.98 2.62 80.66 57.64 56.44
HoliTom [16] 25% 61.2 56.7 58.9 58.4
@@
VisionZip [18] 15% 59.8 54.4 56.1 56.5
PruneVid [27] 10% 59.8 54.5 56.0 56.2
FastVID [44] 10% 56.3 57.3 55.9
LLaVA-Scissor [102] 10% 47.75 3.46 3.26 3.68 3.41 2.90 2.52 80.0 57.5 55.2 57.9
HoliTom [16] 10% 61.2 56.3 56.8 57.3
IV Video-centric Token Compression
@@
To address this, current video LLMs commonly employ a 1 FPS sampling rate to reduce token counts. Furthermore, unlike methods for single images, which often encode both the global image and a series of local patches for detailed feature extraction, video processing often foregoes this detailed frame-level segmentation to keep token numbers manageable. Even with these strategies, the quantity of video tokens remains substantial. During model training and understanding, transformation-based methods, such as the pooling technique used in LLaVA-Video [34], are usually employed to reduce tokens and aid the model’s comprehension of video content.
@@
To fully grasp token compression for video LLMs, it is highly recommended to first review Section III, which details spatial compression methods. Next, we will primarily discuss techniques addressing the temporal domain. Similar to image-centric methods, selected video-centric token compression methods are compared in Table III.
IV-A Transformation-based Video-centric Compression
@@
Like image LLMs, video LLMs use encoders for visual tokens. Consequently, transformation-based video-centric compression methods fundamentally operate on the principles established in Section III-A, with the added capability of performing 3D transformations. A multitude of models showcase cross-modal applicability, performing effectively in both image and video inference tasks. Following the structure of Section III-A, we will now detail transformation-based video-centric compression methods.
IV-A1 2D/3D Pooling
@@
In video LLMs, token pooling is a crucial strategy for managing the high dimensionality of video data. While 2D spatial pooling, as seen in LLaVA-Video [34], can effectively reduce the token count within individual frames, its efficacy alone may be limited for long-duration videos. A growing number of video LLMs, including PLLaVA [3], Video-ChatGPT [90], SlowFast-LLaVA [92], and LongVLM [91], consequently emphasize temporal pooling, which involves downsampling at the frame level.
@@
Notably, PLLaVA’s experiments demonstrate that model performance exhibits greater sensitivity to temporal pooling than to spatial pooling, highlighting its critical role. For extremely long video sequences, LLaMA-VID [40] employs a more aggressive adaptive pooling approach. This method intelligently maintains original resolution for single-image inputs but compresses each video frame to a single token during extended sequence processing, achieving substantial data reduction while aiming to preserve essential information.
@@
This dual focus on spatial and increasingly on temporal pooling underscores their combined importance in enabling efficient processing and comprehensive understanding of video content, particularly as video durations extend. SlowFast-LLaVA [92] incorporates a two-stream SlowFast projector into a LLaVA-style architecture, using a slow pathway to sample fewer, spatially rich frames and a fast pathway to sample more, spatially compressed frames, then concatenates both for the LLM—achieving efficient long-form video understanding with reduced token count while preserving spatiotemporal details.
IV-A2 2D/3D Convolution
@@
Similar to pooling, convolution can also be employed for downsampling video tokens, but it does so in a parameterized manner. Instead of simply aggregating information like pooling, convolution layers learn filters to process and condense spatial and temporal features. VideoLLaMA 2 [10], for instance, thoroughly investigated both 2D and 3D pooling and convolution approaches. Their experiments showed that 3D convolution yielded the best balance of performance and efficiency for video token downsampling. This suggests that learning intricate spatiotemporal relationships through convolutions is more effective for comprehensive video understanding compared to pooling alone.
IV-B Similarity-based Video-centric Compression
@@
Chat-UniVi [94] initially pools each video frame into a single frame-level representation token. It then utilizes DPC-KNN [165, 166] (density peak clustering based on K-nearest neighbors) to amalgamate non-essential frames based on these frame representation tokens. Within each resulting cluster, tokens from multiple frames are further clustered to obtain concise spatiotemporal visual representations. Similarly, FastVID [44] divides video frames solely based on the similarity of their adjacent frame representation tokens. It then employs DPC-KNN within these clustered frames to merge tokens, thereby reducing spatiotemporal redundancy. PruneVid [27] adopts the same frame clustering methodology as Chat-UniVi. The key distinction is that it performs an initial merging of temporally static tokens before executing the spatiotemporal token consolidation. HoliTom [16] argues that relying on a single frame-level representation token for video frame clustering can lead to suboptimal detail capture, and that the preliminary merging of static temporal tokens is disconnected from the original frame clustering method. HoliTom re-conceptualizes temporal redundancy compression as an optimization problem aimed at maximizing the compressible temporal redundant features within all clustered frames, thus addressing temporal compression more holistically. DyCoke [17] groups frames into sets of four, directly performing temporal pruning within each group.
@@
While some methods do not explicitly cluster video frames, FrameFusion [79], for example, acts as a token compression technique for video LLMs. It directly merges temporally redundant tokens exceeding a specific threshold in the shallow layers of the model.
IV-C Attention-based Video-centric Compression
@@
Current attention-based token compression methods in video LLMs and image LLMs share significant similarities. When attention is applied within the encoder to guide token compression, videos are typically treated as a sequence of images fed into an image encoder, making these approaches similar to image-centric token compression. For a more concise discussion of such attention-based methods, please refer to Section III-C.
@@
In contrast, methods employing attention within the decoder process video frames sequentially, concatenating their tokens over time. For longer videos, particularly in the context of streaming video LLMs, windowed attention is commonly used to mitigate computational overhead by focusing on local temporal visual information. However, it’s notable that even these windowed attention-based methods within the decoder often share the same foundational principles as those discussed in Section III-C.
IV-D Query-based Video-centric Compression
IV-D1 Token Distillation
@@
Token distillation in video LLMs commonly relies on specialized adaptor modules, such as the Q-former [1, 81] or Token Turing Machines [104]. These modules typically process video tokens with the learnable query tokens to be attended.
@@
Token Turing Machines (TTMs) [104] maintain a compact external memory of summary tokens, sequentially compressing both new input tokens and memory at each timestep via a Transformer-based read/write mechanism, allowing scalable and efficient processing of long video sequences. BLIP-3-Video [105] introduces an explicit temporal encoder that abstracts hundreds of frame-level visual tokens into as few as 16–32 spatiotemporal tokens using learnable pooling and sequential models, enabling efficient video understanding with limited token usage. LinVT [108] proposes a plug-and-play Linear Video Tokenizer, which linearly aggregates frame-level visual tokens into a compact set of video tokens through spatio-temporal scoring, multi-scale pooling, and text-conditioned aggregation, enabling existing image-LLMs to efficiently process videos and dynamically extract question-relevant information. Long-VMNet [106] accelerates long-form video understanding by using a neural sampler to select discriminative visual tokens from clips and storing them in a fixed-size memory bank for each video; downstream queries are answered by processing only these memory tokens, greatly reducing computational cost while preserving key spatiotemporal information. STORM [107] inserts a Mamba-based [167] temporal encoder between the image encoder and LLM, using spatiotemporal scanning and pooling to inject temporal context into frame tokens and then aggressively compresses tokens by temporal and spatial pooling, enabling efficient long video understanding with minimal token loss. To understand more methods and applications of token distillation in video LLMs, please also refer to Section III-D for a detailed explanation.
IV-D2 Cross-Modal Selection
@@
In video large language models (video LLMs), a query is commonly used to guide the selection of salient frames. In extreme cases, only a handful of frames are relevant to the posed question, allowing the tokens from the vast majority of remaining frames to be discarded. When dealing with an immense number of frames, finding query-relevant information can be akin to searching for a ”needle in a haystack” for the LLM. Query-based token compression methods can pre-filter query-relevant tokens, significantly alleviating the computational burden on the LLM.
@@
LongVU [95] exemplifies this approach. It calculates the relevance of each video frame to the query via cross-modal interaction. This relevance score then dictates a lower compression ratio for key frames, better preserving critical information, all while ensuring the total number of tokens remains within the maximum context length of the LLM.
V Audio-centric Token Compression
@@
The extraction of information from the audio modality can be categorized according to the format of audio representation: (1) continuous sequence modeling: this approach utilizes a pre-trained audio encoder, typically models like Whisper [131] or Conformer [168], to produce continuous audio embeddings; (2) discrete sequence modeling: this method transforms the input audio signal into discrete audio tokens, usually via vector quantization, where continuous audio features are encoded into a learnable codebook. Mainstream methods include HuBERT [169] and EnCodec [170, 171].
@@
V-A Transformation-based audio-centric Compression
Following the categorization in the visual modality, we can classify methods based on their downsampling operations:
V-A1 Token Stacking
Similar to the pixel unshuffle operation in 2D image processing, this approach for audio LLMs token compression involves stacking multiple consecutive tokens along the hidden dimension of the token. This effectively reduces the total number of tokens. Notably, HTS-AT [109], an early example of audio token stacking for classification tasks within audio transformers, utilized 2D pixel-unshuffling on the 2D features extracted from Mel spectrograms to reduce audio tokens. More recent methods such as SLAM-ASR [110], LLaMA-Omni [111], Llama-AVSR [115] and others [172] stack the audio token. Since these token stacking operations alter the hidden dimension, an MLP is typically used to realign the dimension for compatibility with other modalities.
V-A2 Pooling
Another common technique for reducing the number of audio tokens is pooling. Models like Qwen2-audio [113] and Qwen2.5-Omni [5] leverage pooling layers with a stride of 2 to directly decrease the length of the audio representation in a parameter-free manner. This effectively downsamples the audio features, leading to a more compact token sequence. Extending this concept, Llama-MTSK [116] employs a matryoshka-based training approach for flexible token compression. It trains the model with multi-scale audio and video information, achieved by applying average pooling or token stacking at various rates to the initial tokens. This enables Llama-MTSK to dynamically adjust the number of tokens processed during inference, balancing compression and performance within a single model.
V-A3 Temporal Convolution
For audio tokens, 1D convolutions applied across the temporal dimension can reduce the number of tokens. This method simultaneously allows for the alignment of the hidden dimension for subsequent LLM. Approaches like SpeechVerse [112], Baichuan-Audio [114], OSUM [117], and LUCY [118] have employed this technique, often resulting in a downsampled audio representation with an effective sampling rate of 12.5 Hz.
These methods demonstrate how insights from image compression, particularly those involving transformations, can be effectively applied to the audio domain to achieve more efficient token representation for large models.
V-B Similarity-based audio-centric Compression
Similarity-based compression methods aim for each audio token to carry unique information rather than being overly redundant. Similar to the ToMe [21] method used in vision transformers (ViT), A-ToMe [119] inserts a token merge module between the multihead self-attention (MHSA) and feed-forward network (FFN). This module merges adjacent audio tokens that have high cosine similarity.
V-C Attention-based audio-centric Compression

### v5

@@ 3.4.3 Analysis of Similarity Methods
FastV (Chen et al., 2024a) 50% 47.95 3.47 3.36 3.77 3.50 2.99 2.57 81.1 58.0 57.5
DyCoke (Tao et al., 2025a) 50% 47.88 3.47 3.33 3.76 3.51 3.01 2.58 81.1 57.7 57.4 57.5
PLLaVA (Xu et al., 2024) 50% 47.59 3.45 3.36 3.73 3.52 3.00 2.66 81.0 57.7 56.9
LLaVA-Scissor (Sun et al., 2025) 50% 47.89 3.47 3.37 3.76 3.47 3.00 2.65 81.12 57.6 57.4
@@
FastV (Chen et al., 2024a) 35% 47.83 3.46 3.32 3.74 3.47 2.97 2.61 80.5 57.8 56.0 61.3
DyCoke (Tao et al., 2025a) 35% 47.81 3.45 3.31 3.74 3.46 2.98 2.54 80.9 57.7 56.2 61.8
PLLaVA (Xu et al., 2024) 35% 47.23 3.42 3.26 3.70 3.39 2.92 2.59 79.66 56.07 54.26 59.5
VisionZip (Yang et al., 2025c) 25% 60.3 56.5 58.2 57.9
PruneVid (Huang et al., 2025b) 25% 59.9 55.7 57.4 57.4
FastVID (Shen et al., 2025a) 25% 56.3 58.0 56.5
LLaVA-Scissor (Sun et al., 2025) 25% 47.79 3.47 3.33 3.76 3.47 2.98 2.62 80.66 57.64 56.44
HoliTom (Shao et al., 2025) 25% 61.2 56.7 58.9 58.4
@@
VisionZip (Yang et al., 2025c) 15% 59.8 54.4 56.1 56.5
PruneVid (Huang et al., 2025b) 10% 59.8 54.5 56.0 56.2
FastVID (Shen et al., 2025a) 10% 56.3 57.3 55.9
LLaVA-Scissor (Sun et al., 2025) 10% 47.75 3.46 3.26 3.68 3.41 2.90 2.52 80.0 57.5 55.2 57.9
HoliTom (Shao et al., 2025) 10% 61.2 56.3 56.8 57.3
4 Video-centric Token Compression
@@
To address this, current video LLMs commonly employ a 1 FPS sampling rate to reduce token counts. Furthermore, unlike methods for single images, which often encode both the global image and a series of local patches for detailed feature extraction, video processing often foregoes this detailed frame-level segmentation to keep token numbers manageable. Even with these strategies, the quantity of video tokens remains substantial. During model training and understanding, transformation-based methods, such as the pooling technique used in LLaVA-Video (Zhang et al., 2024d), are usually employed to reduce tokens and aid the model’s comprehension of video content.
@@
To fully grasp token compression for video LLMs, it is recommended to first review Section 3, which details spatial compression methods. Next, we will primarily discuss techniques addressing the temporal domain. Similar to image-centric methods, selected video-centric token compression methods are compared in Table 3.
4.1 Transformation-based Video-centric Compression
@@
Like image LLMs, video LLMs use encoders for visual tokens. Consequently, transformation-based video-centric compression methods fundamentally operate on the principles established in Section 3.1, with the added capability of performing 3D transformations. A multitude of models showcase cross-modal applicability, performing effectively in both image and video inference tasks. Following the structure of Section 3.1, we will now detail transformation-based video-centric compression methods.
4.1.1 2D/3D Pooling
@@
In video LLMs, token pooling is a crucial strategy for managing the high dimensionality of video data. While 2D spatial pooling, as seen in LLaVA-Video (Zhang et al., 2024d), can effectively reduce the token count within individual frames, its efficacy alone may be limited for long-duration videos. A growing number of video LLMs, including PLLaVA (Xu et al., 2024), Video-ChatGPT (Maaz et al., 2024), SlowFast-LLaVA (Xu et al., 2025d), and LongVLM (Weng et al., 2024), consequently emphasize temporal pooling, which involves downsampling at the frame level.
@@
Notably, PLLaVA demonstrates that model performance exhibits greater sensitivity to temporal pooling than to spatial pooling, highlighting its critical role. For extremely long video sequences, LLaMA-VID (Li et al., 2024c) employs a more aggressive adaptive pooling approach. This method intelligently maintains original resolution for single-image inputs but compresses each video frame to a single token during extended sequence processing, achieving substantial data reduction while aiming to preserve essential information.
@@
This dual focus on spatial and increasingly on temporal pooling underscores their combined importance in enabling efficient processing and comprehensive understanding of video content, particularly as video durations extend. SlowFast-LLaVA (Xu et al., 2025d) incorporates a two-stream SlowFast projector into a LLaVA-style architecture, using a slow pathway to sample fewer, spatially rich frames and a fast pathway to sample more, spatially compressed frames, then concatenates both for the LLM—achieving efficient long-form video understanding with reduced token count while preserving spatiotemporal details.
Figure 4: Trade-off between Retained Ratio and Performance across Modalities. Left: We visualize changes in token retention and model performance on the VQA2 (Goyal et al., 2017) for image LLMs using each method’s reported setup with LLaVA-1.5-7B (Liu et al., 2023). Right: For video LLMs, we plot the video-token retention ratio and the corresponding performance deltas on the VideoMME benchmark (Fu et al., 2025a), following each method’s reported configuration with LLaVA-OV-7B (Li et al., 2025a). As different methods target distinct compression regimes, we primarily report results at the compression rates specified in their original papers.
4.1.2 2D/3D Convolution
@@
Similar to pooling, convolution can also be employed for downsampling video tokens, but it does so in a parameterized manner. Instead of simply aggregating information like pooling, convolution layers learn filters to process and condense spatial and temporal features. VideoLLaMA 2 (Cheng et al., 2024c), for instance, thoroughly investigated both 2D and 3D pooling and convolution approaches. Their experiments showed that 3D convolution yielded the best balance of performance and efficiency for video token downsampling. This suggests that learning intricate spatiotemporal relationships through convolutions is more effective for comprehensive video understanding compared to pooling alone.
4.2 Similarity-based Video-centric Compression
@@
Chat-UniVi (Jin et al., 2024) initially pools each video frame into a single frame-level representation token. It then utilizes DPC-KNN (Du et al., 2016; Rodriguez and Laio, 2014) (density peak clustering based on K-nearest neighbors) to amalgamate non-essential frames based on these frame representation tokens. Within each resulting cluster, tokens from multiple frames are further clustered to obtain concise spatiotemporal visual representations. Similarly, FastVID (Shen et al., 2025a) divides video frames solely based on the similarity of their adjacent frame representation tokens. It then employs DPC-KNN within these clustered frames to merge tokens, thereby reducing spatiotemporal redundancy. PruneVid (Huang et al., 2025b) adopts the same frame clustering methodology as Chat-UniVi. The key distinction is that it performs an initial merging of temporally static tokens before executing the spatiotemporal token consolidation. HoliTom (Shao et al., 2025) argues that relying on a single frame-level representation token for video frame clustering can lead to suboptimal detail capture, and that the preliminary merging of static temporal tokens is disconnected from the original frame clustering method. HoliTom re-conceptualizes temporal redundancy compression as an optimization problem aimed at maximizing the compressible temporal redundant features within all clustered frames, thus addressing temporal compression more holistically. DyCoke (Tao et al., 2025a) groups frames into sets of four, directly performing temporal pruning within each group.
@@
While some methods do not explicitly cluster video frames, FrameFusion (Fu et al., 2025b), for example, acts as a token compression technique for video LLMs. It directly merges temporally redundant tokens exceeding a specific threshold in the shallow layers of the model.
4.3 Attention-based Video-centric Compression
@@
Current attention-based token compression methods in video LLMs and image LLMs share significant similarities. When attention is applied within the encoder to guide token compression, videos are typically treated as a sequence of images fed into an image encoder, making these approaches similar to image-centric token compression. For a more concise discussion of such attention-based methods, please refer to Section 3.3.
@@
In contrast, methods employing attention within the decoder process video frames sequentially, concatenating their tokens over time. For longer videos, particularly in the context of streaming video LLMs, windowed attention is commonly used to mitigate computational overhead by focusing on local temporal visual information. However, it’s notable that even these windowed attention-based methods within the decoder often share the same foundational principles as those discussed in Section 3.3.
4.4 Query-based Video-centric Compression
4.4.1 Token Distillation
@@
Token distillation in video LLMs commonly relies on specialized adaptor modules, such as the Q-former (Liu et al., 2023; Li et al., 2023c) or Token Turing Machines (Ryoo et al., 2023). These modules typically process video tokens with the learnable query tokens to be attended.
@@
Token Turing Machines (TTMs) (Ryoo et al., 2023) maintain a compact external memory of summary tokens, sequentially compressing both new input tokens and memory at each timestep via a Transformer-based read/write mechanism, allowing scalable and efficient processing of long video sequences. BLIP-3-Video (Ryoo et al., 2024) introduces an explicit temporal encoder that abstracts hundreds of frame-level visual tokens into as few as 16–32 spatiotemporal tokens using learnable pooling and sequential models, enabling efficient video understanding with limited token usage. LinVT (Gao et al., 2024a) proposes a plug-and-play Linear Video Tokenizer, which linearly aggregates frame-level visual tokens into a compact set of video tokens through spatio-temporal scoring, multi-scale pooling, and text-conditioned aggregation, enabling existing image-LLMs to efficiently process videos and dynamically extract question-relevant information. Long-VMNet (Gurukar and Kadav, 2025) accelerates long-form video understanding by using a neural sampler to select discriminative visual tokens from clips and storing them in a fixed-size memory bank for each video; downstream queries are answered by processing only these memory tokens, greatly reducing computational cost while preserving key spatiotemporal information. STORM (Jiang et al., 2025) inserts a Mamba-based (Gu and Dao, 2024a) temporal encoder between the image encoder and LLM, using spatiotemporal scanning and pooling to inject temporal context into frame tokens and then aggressively compresses tokens by temporal and spatial pooling, enabling efficient long video understanding with minimal token loss. To understand more methods and applications of token distillation in video LLMs, please also refer to Section 3.4 for a detailed explanation.
4.4.2 Cross-Modal Selection
@@
In video large language models (video LLMs), a query is commonly used to guide the selection of salient frames. In extreme cases, only a handful of frames are relevant to the posed question, allowing the tokens from the vast majority of remaining frames to be discarded. When dealing with an immense number of frames, finding query-relevant information can be akin to searching for a "needle in a haystack" for the LLM. Query-based token compression methods can pre-filter query-relevant tokens, significantly alleviating the computational burden on the LLM.
@@
LongVU (Shen et al., 2025b) exemplifies this approach. It calculates the relevance of each video frame to the query via cross-modal interaction. This relevance score then dictates a lower compression ratio for key frames, better preserving critical information, all while ensuring the total number of tokens remains within the maximum context length of the LLM.
5 Audio-centric Token Compression
@@
The extraction of information from the audio modality can be categorized according to the format of audio representation: (1) continuous sequence modeling: this approach utilizes a pre-trained audio encoder, typically models like Whisper (Radford et al., 2023) or Conformer (Gulati et al., 2020), to produce continuous audio embeddings; (2) discrete sequence modeling: this method transforms the input audio signal into discrete audio tokens, usually via vector quantization, where continuous audio features are encoded into a learnable codebook. Mainstream methods include HuBERT (Hsu et al., 2021) and EnCodec (Défossez et al., 2022; Zeghidour et al., 2021).
@@
5.1 Transformation-based audio-centric Compression
Following the visual modality’s categories, we can classify methods based on their downsampling operations:
5.1.1 Token Stacking
Similar to the pixel unshuffle operation in 2D image processing, this approach for audio LLMs token compression involves stacking multiple consecutive tokens along the hidden dimension of the token. This effectively reduces the total number of tokens. Notably, HTS-AT (Chen et al., 2022), an early example of audio token stacking for classification tasks within audio transformers, utilized 2D pixel-unshuffling on the 2D features extracted from Mel spectrograms to reduce audio tokens. More recent methods such as SLAM-ASR (Ma et al., 2024c), LLaMA-Omni (Fang et al., 2024), Llama-AVSR (Cappellazzo et al., 2025a) and others (Fathullah et al., 2024) stack the audio token. Since these token stacking operations alter the hidden dimension, an MLP is typically used to realign the dimension for compatibility with other modalities.
5.1.2 Pooling
Another common technique for reducing the number of audio tokens is pooling. Models like Qwen2-audio (Chu et al., 2024b) and Qwen2.5-Omni (Xu et al., 2025b) leverage pooling layers with a stride of 2 to directly decrease the length of the audio representation in a parameter-free manner. This effectively downsamples the audio features, leading to a more compact token sequence. Extending this concept, Llama-MTSK (Cappellazzo et al., 2025b) employs a matryoshka-based training approach for flexible token compression. It trains the model with multi-scale audio and video information by applying average pooling or token stacking at different rates to the initial tokens. This enables Llama-MTSK to dynamically adjust the number of tokens processed during inference, balancing compression and performance within a single model.
5.1.3 Temporal Convolution
For audio tokens, 1D convolutions applied across the temporal dimension can reduce the number of tokens. This method simultaneously allows for the alignment of the hidden dimension for subsequent LLM. Approaches like SpeechVerse (Das et al., 2024), Baichuan-Audio (Li et al., 2025b), OSUM (Geng et al., 2025), and LUCY (Gao et al., 2025) have employed this technique, often resulting in a downsampled audio representation with an effective sampling rate of 12.5 Hz.
These methods demonstrate how insights from image compression, particularly involving transformations, can be effectively applied to the audio domain to achieve more efficient token representation for large models.
5.2 Similarity-based audio-centric Compression
Similarity-based compression methods aim for each audio token to carry unique information rather than being overly redundant. Similar to the ToMe (Bolya et al., 2022) method used in vision transformers (ViT), A-ToMe (Li et al., 2023f) inserts a token merge module between the multihead self-attention (MHSA) and feed-forward network (FFN). This module merges adjacent audio tokens that have high cosine similarity.
5.3 Attention-based audio-centric Compression